16. Suppose the supply side of an economy is characterized by:

A production function: Q = 3LK; where Q = output, L = labor, and

K= capital

Labor demand: LD = 10 - 2W/P , where W = nominal wage, P = price

Labor supply: LS = 4P/W ; and

Capital stock is fixed at K = 4.

(a) Derive the aggregate supply curve.

(b) Which cases (intermediate Keynessian, extreme Keynesian, classical) does the supply curve represent?

(c) How would your answers to (a) and (b) change if nominal wage

was fixed at 3?


1
Expert's answer
2022-01-17T12:46:29-0500

"Solution"

"4p^2=w(10-2w)\\
4p^2=10w-2w^2\\
Ls=4p^2-2w^2+10w\\
Ls=8p-4w+10"


B)extreme Keynesian

C)the labor supplied will decrease